What deposit made at the beginning of each quarter will accumulate to $36 000.00 at 4% compounded quarterly at the end of eight years? Select one: a. $950.65 b. $317.94 c. $960.15 d. $314.78 e. $318.99
Answer #1

Answer is $950.65

Desired Sum = $36,000
Time Period = 8 years or 32 quarters

Annual Interest Rate = 4.00%
Quarterly Interest Rate = 1.00%

Let deposit at the beginning of each quarter be $x

$x*1.01^32 + $x*1.01^31 + … + $x*1.01^2 + $x*1.01 = $36,000
$x * 1.01 * (1.01^32 - 1) / 0.01 = $36,000
$x * 37.8690085 = $36,000
$x = $950.65

Quarterly Deposit = $950.65